The demand function for a product is: Q = 1,000-10P. and its supply function is: Qs = 400 + 5P.

The equilibrium price of the good is equal to ____

The equilibrium quantity of the good is equal to ______
Answer #1

The equilibrium price of the good is equal to $40.

The equilibrium quantity of the good is equal to 600.

Explanation:

For equilibrium:

Q​​​​​​d = Q​​​​​​s

1,000 - 10P = 400 + 5P

15P = 600

P = 600 / 15 = $40

Q = 1,000 - 10P = 1,000 - 10(40) = 600